OpenStudy (anonymous):

If (x^2 + ax - 10) ÷ (x + 2) = (x - 5), then a =

OpenStudy (anonymous):

A. -7 B. -3 C. 7 D. 3

Parth (parthkohli):

Then, \(x^2 + ax - 10 = (x - 5)( x+ 2)\) Do you get it till here?

OpenStudy (anonymous):

kinda.....

Parth (parthkohli):

Yeah. I just multiplied both sides by \(x + 2\).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yeah i get it

Parth (parthkohli):

Great! Can you simplify \((x - 5)(x + 2)\)?

OpenStudy (anonymous):

Middle term will be negative because here the factors are -5 and 2.. So "a" will be negative..

OpenStudy (anonymous):

@ParthKohli do you mean setting them equal to 0?

OpenStudy (anonymous):

So your answer can be either A or B.. You have to check now the sum: -5 + 2 = ??

Parth (parthkohli):

Nope. \(x^2 + a x - 10 = (x + 5)( x - 2)\) You have to simplify the right hand side by multiplying x + 5 and x - 2.

OpenStudy (anonymous):

oh ok so its x^2+3x-10?

OpenStudy (anonymous):

@ParthKohli are you there?

Parth (parthkohli):

Yep :)

Parth (parthkohli):

Sorry for the late reply, I was eating my dinner.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x + 5)(x - 2) = x(x - 2) + 5(x - 2) = x^2 - 2x + 5x - 10 = x^2 + 3 x - 10 }\)

Parth (parthkohli):

So what is \(a\)?